### v1

# FSFSplatter: Build Surface and Novel Views with Sparse-Views within 3min

Abstract
Gaussian Splatting has become a leading reconstruction technique, known for its high-quality novel view synthesis and detailed reconstruction. However, most existing methods require dense, calibrated views. Reconstructing from free sparse images often leads to poor surface due to limited overlap and overfitting. We introduce FSFSplatter, a new approach for fast surface reconstruction from free sparse images. Our method integrates end-to-end dense Gaussian initialization, camera parameter estimation, and geometry-enhanced scene optimization. Specifically, FSFSplatter employs a large Transformer to encode multi-view images and generates a dense and geometrically consistent Gaussian scene initialization via a self-splitting Gaussian head. It eliminates local floaters through contribution-based pruning and mitigates overfitting to limited views by leveraging depth and multi-view feature supervision with differentiable camera parameters during rapid optimization. FSFSplatter outperforms current state-of-the-art methods on widely used DTU and Replica.
1 Introduction
Surface reconstruction and novel view synthesis (NVS) from unconstrained multi-view RGB images are long-standing tasks in autonomous driving[1], computer graphics[2], and virtual reality[3]. As typical methods require dense, calibrated images, accurate reconstruction and NVS from sparse, free inputs remain particularly challenging, yet crucial for scenarios with only limited RGB captures.
Traditional approaches often decompose the problem into separate stages of camera pose estimation, dense reconstruction, and surface extraction, typically by converting dense point clouds into surfaces via signed distance functions (SDFs). This pipeline, however, introduces errors at each stage, leading to irreversible error accumulation.
Recent studies have explored end-to-end models for unconstrained image reconstruction. For example, Dust3R[4] learns a robust image-to-PointMap mapping from image pairs, with several variants[5, 6]. Nevertheless, its network architecture is incapable of processing multiple views in a single forward pass, resulting in accumulated inconsistencies. VGGT[7] employs alternating global and local attention mechanisms, enabling 3D reconstruction from an arbitrary number of unconstrained images.
The sparsity of point clouds, however, makes them a suboptimal representation for many downstream applications (e.g., NVS). Recent breakthroughs in Neural Radiance Fields (NeRF)[8] and differentiable rendering[9] have demonstrated unprecedented quality in multi-view reconstruction. However, most of these methods struggle to generalize to unconstrained sparse-view settings, as insufficient image overlap often leads to the failure of Structure-from-Motion (SfM), and the inherent ambiguity in dense reconstruction arises from view extrapolation and occlusion under sparse views. Methods like PF-LRM[10, 11] employ a Transformer to map RGB images into a NeRF representation. Nevertheless, the surface bias of NeRF[12] and the separate point cloud head for camera pose estimation limit the final reconstruction quality. In contrast, Gaussian Splatting (GS) scenes, due to their explicit scene representation and unbiased geometry, have shown advantages in surface reconstruction[13, 2]. FreeSplatter[14] constructs a GS scene in a feed-forward manner from a single, unconstrained RGB image, and subsequently estimates camera parameters and aligns multiple views using an optimizer.
We employ a large transformer as the backbone network, which regresses scale-consistent camera parameters, depth maps, and a Gaussian scene via independent prediction heads. To predict dense and geometrically accurate Gaussian primitives, we specifically design a geometry and global feature-based Gaussian self-splitting decoder. A semi-dense Gaussian scene is first obtained by fusing the pixel-aligned depth map and DPT-decoded features. Subsequently, an Encoder-Decoder architecture is utilized to compute offset maps and perform Gaussian self-splitting, yielding the dense Gaussian scene. To enhance the overall geometric stability and eliminate local floaters, we apply a contribution-based pruning strategy to obtain the initial scene. Furthermore, to address the issue of local view overfitting during the optimization process under sparse inputs, we propose an optimization scheme supervised by monocular depth and multi-view features, which significantly improves both surface reconstruction and NVS quality within a short time.
Our main contributions can be summarized as follows:
  1. 1.
    We propose FSFSplatter, a fast surface reconstruction algorithm from free-sparse images. Compared to prior works, our method achieves accurate reconstruction results both with and without per-scene optimization.
  2. 2.
    The Gaussian head performs Gaussian self-splitting based on high-dimensional features, coupled with contribution-based pruning, enabling precise initialization of the dense Gaussian scene. Furthermore, we introduce a geometry-enhanced scene optimization supervised by monocular depth and multi-view features, which effectively addresses the common issue of local view overfitting in sparse-view optimization.
  3. 3.
    We demonstrate state-of-the-art (SOTA) performance on widely-used DTU and Replica datasets upon surface reconstruction and NVS.
2 Related Works
2.1 Surface Reconstruction
Surface reconstruction from RGB images is a fundamental problem in computer vision. Traditional methods typically represent scenes as point clouds [15], depth maps [16], or voxels [17]. In recent years, with the advancement of computational hardware and neural network technologies, implicit scene representations have been widely adopted. NeRF and volume rendering [8] have been applied to NVS and 3D reconstruction, undergoing rapid development. To address the surface bias inherent in volume rendering, NeuS [12, 18] proposed an unbiased NVS-based surface reconstruction method, achieving accurate surface reconstruction based on an implicit signed distance function (SDF) and demonstrating the advantages of NVS-based surface reconstruction in detailed regions.
3D Gaussian Splatting (3DGS) represents a scene using a set of explicit Gaussian primitives [9], enabling fast and high-quality NVS and reconstruction. However, 3DGS lacks expressive capacity for scene geometry. Subsequent work has achieved accurate surface reconstruction by constraining the shapes of Gaussian primitives to align with the surface [2, 13]. Through depth map rendering and fusion, fast and accurate geometric surface reconstruction can be realized. Furthermore, mapping algorithms from colored point clouds to Gaussians have further enhanced the reconstruction quality [19, 20].
2.2 Pose-free Reconstruciton
However, most of the aforementioned methods require dense camera views with known parameters. Under sparse views, Structure-from-Motion (SfM) often fails due to insufficient image overlap. Even with constrained camera parameters, errors in the camera parameters and sparsity can gradually lead to overfitting to individual views, resulting in erroneous geometric surfaces.
In traditional pipelines, image-based 3D reconstruction involves separate stages such as feature extraction, feature matching, bundle adjustment (BA), and plane sweeping. Colmap [15, 21], introduced in 2016, has become a benchmark method for sparse and dense reconstruction due to its exceptional robustness and performance.
However, both camera pose estimation from sparse views and NVS-based dense reconstruction methods suffer from unavoidable ambiguities caused by extrapolation and occlusion [22]. This leads the scene to overfit the sparse color observations, forming incorrect geometry that severely degrades the quality of surface reconstruction. Some approaches address this by leveraging large-scale dataset pre-training for better generalization to other datasets, but often require several days of training [22]. Methods focusing on per-scene optimization introduce additional supervision to enhance geometric stability during the optimization process [23, 24]. These methods typically incorporate monocular depth estimation [25] as a geometric prior to stabilize the optimization.
With the growing recognition of the potential of large Transformer models, pose estimation and 3D reconstruction are increasingly being implemented in an end-to-end manner, thereby mitigating the issue of cumulative errors inherent in the aforementioned multi-stage pipelines. Dust3R [4] and its subsequent developments [5, 6] achieve accurate camera parameter estimation and PointMap prediction from just two images. VGGT facilitates joint camera parameter estimation and 3D reconstruction from multiple images in a single forward pass by leveraging an alternating mechanism of global and local attention [7].
Beyond scene representations based on point clouds and depth maps, several works build upon NeRF and Gaussian Splatting (GS) representations. PF-LRM [10] maps unconstrained multi-view RGB images to a NeRF representation via a transformer, employing separate branches to predict point clouds and camera poses, which introduces additional training costs and challenges for surface reconstruction. FreeSplatter [14] maps unconstrained single views to a Gaussian scene representation and performs multi-view camera pose estimation and scene alignment based on optimization, treating scene generation and camera pose estimation as independent steps.
3 Method
We propose a fast reconstruction method based on free-sparse RGB images that enables robust 3D surface reconstruction from only three input images within 3 minutes.
Our framework consists of two main components: an end-to-end differentiable dense Gaussian initialization and camera parameter estimation, and a geometry-aware robust Gaussian scene optimization.
3.1 Preliminaries
3.1.1 Gaussian Splatting
Based on Gaussian Splatting [9], which represents a scene as a collection of Gaussian primitives, 2DGS further flattens these 3D Gaussians into 2D primitives [2], enabling precise surface reconstruction by closely adhering the Gaussians to the actual surfaces. Each Gaussian primitive is defined by its position $\mu^{3}$, rotation $R^{4}$, scale $S^{2}$, opacity $O^{1}$, and spherical harmonics (SH) coefficients $C^{48}$ for representing color.
The splatting of Gaussian primitives from the Gaussian coordinate system to the pixel coordinate system is achieved via a homogeneous transformation matrix $\mathbf{H}$, as shown in Equation 1.
$\mathbf{u}$ $=\mathbf{W}\mathbf{H}(s_{1},s_{2},1,1)^{\top}$ (1)
Multiple Gaussian primitives are composited via $\alpha$-blending, as shown in Equation 2. Here, $\mathcal{G}$ denotes the set of primitives influencing the pixel, and $\alpha$ is computed based on the Gaussian’s covariance matrix $\boldsymbol{\Sigma}$, its center coordinates $\boldsymbol{p}$, and the projected pixel coordinates $\boldsymbol{\mu}$.
$\mathbf{C}$ $=\sum_{i\in\mathcal{G}}\mathbf{c}_{i}\alpha_{i}\prod_{i-1}^{n}(1-\alpha_{n})$ (2)
$where\ \ \alpha_{n}$ $=\sigma_{n}\cdot\exp\left(-\frac{1}{2}\left(\boldsymbol{p}-\boldsymbol{\mu}_{n}\right)^{T}\boldsymbol{\Sigma}_{n}^{-1}\left(\boldsymbol{p}-\boldsymbol{\mu}_{n}\right)\right)$
3.1.2 Visual Geometry Grounded Transformer
To address the issue of error accumulation caused by multiple stages in traditional 3D reconstruction pipelines, VGGT [7] employs a large feed-forward neural network to directly infer all key attributes from multiple input views. These attributes include camera parameters $\mathbf{g}_{i}$, depth maps $D_{i}$, PointMaps $P_{i}$, and point trajectories $T_{i}$. The backbone network incorporates alternating global and local attention mechanisms. Utilizing a shared backbone with separate, dedicated prediction heads enables the simultaneous estimation of all required 3D attributes.
3.2 Problem definition
The input consists of three RGB images $I_{i}\in\mathbb{R}^{3\times 3\times H\times W}$ observing the same scene. Our algorithm $\mathcal{F}$ maps these images to a dense Gaussian representation of the scene, enabling accurate and fast novel view synthesis (NVS) and surface reconstruction through rasterization, as shown in Equation 3.
$\mathcal{F}(\left(I_{i}\right)^{3})=(\mathcal{G},(\mathcal{C}_{i})^{3})$ (3)
where $\mathcal{G}$ denotes the Gaussian scene. This work characterizes the Gaussian scene as a collection of 2D Gaussians within the 3D space, achieving precise alignment with surfaces and normals by flattening the 3D Gaussians [2]. Each Gaussian primitive comprises coordinates $X^{3}$, rotation $R^{3}$, scale $S^{2}$, opacity $O^{1}$, and spherical harmonic coefficients $\text{SH}^{48}$. $(\mathcal{C}_{i})^{3}$ represents the camera parameters (including both intrinsic and extrinsic parameters).
3.3 End-to-end dense Gaussian initialization and camera parameter estimation
Figure 1: The end-to-end dense Gaussian initialization and camera parameter estimation take sparse images as input. Image tokens are processed via a DINO Encoder and alternating attention mechanisms. A Camera Head and a DPT Head perform back-projection and depth estimation, yielding a semi-dense Gaussian scene. This scene is partitioned into multiple patches, which are then fed into a densification MLP to produce a dense Gaussian scene serving as the initialization.
Conventional 3D reconstruction pipelines are typically divided into stages such as feature extraction, feature matching, triangulation, and dense reconstruction. Each stage introduces errors that inevitably accumulate throughout the process.
Inspired by recent advances in 3D depth estimation [7, 4], we adopt a large Transformer architecture as the backbone for our end-to-end dense Gaussian initialization and camera parameter estimation framework. As illustrated in Figure 1, images are first encoded into tokens $(T_{i})^{3}$ using DINOv2 [26]. These tokens are then processed through the main network structure, which employs alternating global and local attention mechanisms.
Previous methods have demonstrated remarkable performance by incorporating decoder heads for camera parameters, depth estimation (DPT), and PointMap prediction [7]. To improve the initialization quality of the Gaussian Splatting (GS) scene, we propose a novel self-splitting-based patch initialization strategy for 2DGS scenes.
Figure 2: Explanation of dense Gaussian initialization. By incorporating explicit normal estimation and an MLP-based implicit Gaussian densification process, the quality of both novel NVS and surface reconstruction is significantly improved.
Based on the outputs from the camera head and the DPT head, we can rapidly perform back-projection to obtain pixel-aligned 3D point clouds. Concurrently, we design an independent DPT back-projection module that utilizes positional encoding to back-project the encoded tokens into pixel-aligned high-dimensional features $\mathcal{F}_{P}$. To address scale variations across different scenes, we first apply normalization to the point cloud. We employ an encoder-decoder architecture to map the aforementioned information into dense Gaussian primitive attributes. Specifically, PointMLP [27] is adopted to encode the geometric information of the point cloud, yielding high-dimensional geometric features $\mathcal{F}^{c}$. Subsequently, $\mathcal{F}_{P}$ is concatenated channel-wise with explicit Gaussian attributes—including $\text{SH}^{48}_{P}$, $R^{4}_{P}$, $S^{2}_{P}$, and $\mathcal{F}_{P}$—to form the encoded representation. We then design mutually independent decoders for the attribute variations $\Delta\mathcal{G}$ of all Gaussian primitives. These decoders predict the densification tendency of the original semi-dense Gaussian scene, as visualized in Fig. 2. Each Gaussian primitive splits into $N$ new primitives based on its predicted variation, as formulated in Equation 4.
$\Delta\mathcal{G}_{N\cdot P}=\{\Delta X^{3}_{N\cdot P},\Delta R^{4}_{N\cdot P},\Delta S^{2}_{N\cdot P},\Delta O^{1}_{N\cdot P},$ $\Delta\text{SH}^{48}_{N\cdot P}\}=D(\text{Cat}(\mathcal{F}_{P},\text{SH}^{48}_{P},R^{4}_{P},S^{2}_{P},\mathcal{F}_{P}))$ (4)
$\text{Densify}(\mathcal{G}_{P})$ $=\mathcal{G}_{P}+\Delta\mathcal{G}_{N\cdot P}$
Drawing on advanced point cloud processing techniques [28, 20], we adopt point cloud patches as the input units to our network to enhance its generalization capability across diverse scenes. After obtaining the back-projected 3D point cloud via the DPT head and camera head, and applying normalization, we employ K-Nearest Neighbors (KNN) to partition the point cloud into randomly sampled local patches. These patches are divided into multiple small batches, each processed through an encoder-decoder architecture to generate densified Gaussian primitives. Finally, after de-normalization, the processed batches are merged to form the complete Gaussian scene, as expressed in Equation 5. Here, $\mathbf{N}$ and $\mathbf{N}^{\prime}$ denote the normalization and de-normalization operations, respectively; $\mathbb{B}$ and $\mathbb{B}^{\prime}$ represent batch partitioning and merging; and $\mathbb{D}$ signifies the aforementioned densification process.
$\mathcal{G}=\mathbf{N}^{\prime}\{\mathbb{B}^{\prime}[\mathbb{B}(\mathbf{N}(\mathcal{G}_{init}))+\mathbb{D}\left(\mathbb{B}\left(\mathbf{N}\left(\mathcal{G}_{init}\right)\right)\right)]\}$ (5)
3.3.1 Training strategy
Our entire pipeline is fully end-to-end differentiable, enabling comprehensive training based on $\mathcal{L}_{\text{RGB-D}}$ and $\mathcal{L}_{\text{Cam}}$. Here, $\mathcal{L}_{\text{RGB-D}}$ comprises an RGB-based loss combining $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$, along with a depth-based loss $\mathcal{L}_{\text{D}}$. The camera parameters are defined by the intrinsic matrix $\mathbf{F}\in\mathbb{R}^{1}$ and the extrinsic parameters $\mathbf{T}=[\mathbf{t}\in\mathbb{R}^{3},\mathbf{q}\in\mathbb{R}^{4}]$, forming the camera loss $\mathcal{L}_{\text{Cam}}=\mathcal{L}_{\mathbf{F}}+\mathcal{L}_{\mathbf{t}}$.
A straightforward approach involves training all parameters of the entire pipeline directly. However, in practice, since both the input and the final pipeline output are RGB images, this often leads to overfitting on the novel view synthesis (NVS) of local views during training. This overfitting can result in erroneous geometry and degrade the quality of surface reconstruction [14]. To mitigate this, we adopt a geometry-progressive training strategy. We initialize our network backbone with pre-trained weights from VGGT-1B [7]. On this foundation, we remove the PointMap Head and Tracking Head, and proceed to train the DPT Head, Camera Head, and the backbone network. During this stage, the losses $\mathcal{L}_{\text{D}}$ and $\mathcal{L}_{\text{Cam}}$ are enabled to ensure geometric stability, maintaining consistency with the training methodology of VGGT.
To achieve a high-quality initialization of the Gaussian scene and prevent overfitting commonly associated with end-to-end NVS training, we subsequently freeze the DPT Head, Camera Head, and the backbone network. This freezing preserves geometric stability during the subsequent training phase. The DPT unproject Head and the densification MLP are then trained independently. In this stage, KNN Patch partitioning is disabled, allowing for global densification and estimation of Gaussian primitives across the entire scene, which serves as a form of pre-training. The losses $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ are enabled during this phase.
Finally, we unfreeze all parameters of the pipeline, enabling the backbone network to learn the mapping from RGB images to Gaussian attributes directly. The KNN Patch operation is reinstated during the densification process to enhance generalization. The complete set of loss functions is activated throughout this final training stage.
3.4 Geometrically-enhanced scene optimization
Figure 3: After initialization, a contribution-based Gaussian pruning process is first applied to eliminate erroneous floaters present in the dense Gaussians. Both the Gaussian primitives and the camera poses are treated as differentiable variables. Multi-view RGB, depth, and normal maps are rendered through rasterization. Various loss functions are constructed based on monocular depth estimation and multi-view high-dimensional feature extraction, with subsequent backpropagation.
3.4.1 Differentiable Gaussian scene construction
2D/3DGS jointly optimizes NVS and geometry through RGB-based losses [2, 9]. However, due to sparse viewpoint coverage and unknown camera parameters, direct optimization using RGB-based losses introduces inevitable multi-view ambiguity [22]. This often leads to color overfitting on sparse views and results in catastrophic geometry degradation. To address this, we propose a geometry-enhanced optimization strategy for Gaussian scenes.
The dense Gaussian scene initialization yields Gaussian primitives that are more densely distributed than pixels. While this provides a strong prior, it also introduces numerous occluded or invisible primitives. During optimization, these non-contributing primitives do not receive gradient updates through backpropagation, making them impossible to prune simply by filtering low-opacity elements. Such primitives are geometrically meaningless and often detrimental to surface reconstruction. Therefore, we perform contribution-based Gaussian pruning prior to optimization.
Referring to Equation 2, the contribution $\mathbf{C}_{n}$ of a Gaussian primitive $\mathcal{G}n$ can be quantified by its weight in the $\alpha$-blending process: $\alpha_{i}\prod{i-1}^{n}(1-\alpha_{n})$, as shown in Equation 6. Here, $\mathcal{P}_{n}$ denotes the set of pixels influenced by the Gaussian primitive $\mathcal{G}_{n}$. To prevent disproportionately large contributions from Gaussians covering extensive areas, we normalize the contribution by the number of occupied pixels $|\mathcal{P}_{n}|$ during accumulation. After computing contributions for all primitives, we sort them and prune those with low contribution or excessively low opacity, as shown in Figure 4.
$\mathbf{C}_{n}=\sum_{k=1}^{3}\frac{1}{|\mathcal{P}_{n}|}\sum_{p\in\mathcal{P}_{n}}(\alpha_{i}^{p}\prod_{i-1}^{n}(1-\alpha_{n}^{p}))$ (6)
Figure 4: Contribution-based Gaussian pruning for scene initialization
2DGS employs a CUDA-based differentiable rasterizer that explicitly computes the gradients of Gaussian primitives. However, the original rasterizer does not support the backward propagation of camera parameters, which is detrimental for pose-free 3D reconstruction, as it tends to amplify subtle multi-view inconsistencies during the optimization process. To address this limitation, we incorporate differentiable camera poses into the differentiable Gaussian primitives. Following related works such as GS-SLAM [29], we consistently position the camera pose at the origin of the coordinate system. Prior to rasterization, we apply the inverse transformation of the camera pose $T^{\text{cam}}_{k}$ to all Gaussian primitives, thereby achieving the same view synthesis effect as in the original framework. Throughout this process, $T^{\text{cam}}_{k}$ remains differentiable. Consequently, any loss based on the rasterized image will propagate back to the camera pose before being passed to the Gaussian primitives during backpropagation. Owing to the sparsity of viewpoints, our approach enables rapid and mutually independent pose optimization across different views.
3.4.2 Loss Function
During the optimization process, our loss function comprises several components: the RGB-based $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$, the normal map-based $\mathcal{L}_{n}$, the losses derived from the rendered depth map and monocular depth estimation-namely $\mathcal{L}_{\text{rank}}$, $\mathcal{L}_{\text{smooth}}$, and $\mathcal{L}_{d}$-as well as $\mathcal{L}_{\text{MVS}}$, which is based on RGB feature extraction and rendered depth back-projection, as illustrated in Figure3. Among these, $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ are consistent with 3DGS [9]; $\mathcal{L}_{d}$ and $\mathcal{L}_{n}$ are are consistent with 2DGS [2]. Furthermore, owing to the dense initialization of the Gaussian scene, we are able to apply all loss supervision from the beginning, without the need for a gradual progression starting solely from the RGB Loss.
We employ Marigold, a Vision Transformer-based monocular depth estimation network, as the geometric prior [30]. The generalized monocular depth estimation $D_{\text{est}}$ ensures the overall stability of the scene during optimization. The 2D Gaussian primitives achieve accurate depth map rendering by closely adhering to the actual surfaces. To address the scale ambiguity inherent in monocular depth estimation, we utilize a depth ranking loss to supervise the rasterized depth $D_{\text{re}}$, as shown in Equation 7. Specifically, pixel coordinates $p_{1},p_{2}$ are sampled from neighboring patches $\mathcal{P}$ using the same random strategy $\mathcal{R}$ applied to both $D_{\text{est}}$ and $D_{\text{re}}$, in order to compute the depth difference. Here, sgn denotes the sign function, and $\sigma$ represents the ReLU function.
$\mathcal{L}_{\text{rank}}=\sum_{p_{1},p_{2}\in\mathcal{P}}\sigma\left(\operatorname{sgn}(\mathcal{S}(D_{\text{pre}},p_{1},p_{2})\cdot\mathcal{S}(D_{\text{re}},p_{1},p_{2}))+m\right)$ (7)
$where\ \ \ \mathcal{S}(D,p_{1},p_{2})=D(p_{1})-D(p_{2})$
Concurrently, we enforce a depth smoothness loss $\mathcal{L}_{\text{smooth}}$ in regions where the monocular depth estimation does not exhibit rapid variations, to preserve the smoothness of the synthesized surfaces. As shown in Equation 8, where $n_{1}$ and $n_{2}$ are small thresholds used to identify edges in the monocular depth estimate.
$\mathcal{L}_{\text{smooth}}=\sum_{p_{1},p_{2}\in\mathcal{P}}\text{sgn}(|D_{\text{pre}}(p_{1})-D_{\text{pre}}(p_{2})|-n_{1})\cdot\sigma((|D_{\text{re}}(p_{1})-D_{\text{re}}(p_{2})|-n_{2})$ (8)
Loss based on monocular depth estimation effectively preserves the overall geometric structure of the scene, yet often lacks sufficient supervision for fine details. While direct RGB supervision of novel views could theoretically provide detailed guidance, inconsistencies in multi-view illumination and camera capture variations introduce color discrepancies across different views [31]. These discrepancies are further exacerbated in free-sparse views, leading to inconsistencies during training.
To address this issue, we employ a multi-view feature alignment loss based on a U-Net architecture, which supervises high-dimensional features derived from both RGB data and rendered depth back-projection. Specifically, we extract high-dimensional features from the input RGB images using a pre-trained U-Net [32]. These features, along with the rendered depth map, are back-projected to reference views. The feature difference between the current view and reference views is then computed as shown in Equation 9.
$\mathcal{L}_{\text{MVS}}=\sum_{r}\left|1-\cos\left[\mathcal{R_{r}}\left(\mathbf{U}^{(c)}\left(I_{r}\right)\right),\mathbf{U}^{(c)}\left(I\right)\right]\right|$ (9)
During training, to accelerate the optimization process, we control the downsampling resolution based on image frequency [33]. Unlike sparse point clouds initialized by Colmap, the dense Gaussian initialization allows the initial training downsampling rate to be set to 2. Owing to the sub-pixel level density of the initial Gaussians, we do not perform gradient-based densification during optimization. Instead, we only execute Gaussian pruning based on opacity and contribution to eliminate low contributing Gaussian primitives and error floaters.
4 Experiments
4.1 Settings
Training
Our Gaussian initialization parameter $K$ is set to 2, with an image resolution of $512\times 512$ and a single patch size of 2048. All our experiments are conducted on a server equipped with an Intel Platinum 8470Q processor and an Nvidia RTX5090 GPU, requiring 9 hours to complete 120 training epochs. The training dataset consists of BlendMVS [34], DTU [35], Virtual KITTI [36], and Replica [37], encompassing extensive RGB-D data and camera parameters from both synthetic and real-captured sources.
Metrics
We employ widely adopted metrics including PSNR, SSIM[38], and LPIPS[39] to evaluate NVS quality, while utilizing Chamfer Distance (CD)[35] to evaluate surface reconstruction quality.
Baselines
We evaluate the reconstruction results on base methods: 3D Gaussian Splatting 3DGS [9]; surface reconstruction method 2DGS [2]. Pose-free methods: RGB based pose-free reconstruction method CF-3DGS [40]; free-sparse-view feed-forward Gaussian Splatting generation method FreeSplatter [14]; dense reconstruction method based on large Transformer VGGT. Sparse-view surface reconstruction methods: sparse-view surface reconstruction method with additional regularization FatesGS; surface reconstruction method incorporating unbiased depth estimation PSGR.
Among these algorithms, 3DGS, 2DGS, FatesGS, and PSGR, as classical GS-based per-scene optimization pipelines, require camera parameters as prior knowledge. We therefore employ VGGT to estimate the camera parameters and subsequently evaluate the performance of these methods under both ground truth camera parameters and estimated camera parameters. Furthermore, we conduct comparative evaluations under two training configurations: one using our fixed iteration count of 1000, and the other using their default iteration settings.
Datasets
We evaluate our method on DTU datasets [35] and Replica datasets [37]. DTU datasets have been extensively used in prior surface reconstruction works. We evaluate on all 15 scenes from DTU, where each scene contains 49 or 64 images at a resolution of $1600\times 1200$. Following previous works [23, 41], we select views 23, 24, and 33 as input and downsample the images to $800\times 600$. Since DTU datasets primarily consist of object-level scenes, we further validate our algorithm on Replica datasets to assess its generalization capability and performance on scene-level scenes. We select the first five scenes and randomly choose three input views at a resolution of $1200\times 680$.
4.2 Evaluation of Surface Reconstruction
We first evaluate the surface reconstruction accuracy (CD) under different settings across all baselines. As shown in Table1, Table2, and Figure5, our method demonstrates significant advantages in CD error even when compared with per-scene optimization methods using their default iteration numbers, and remains competitive with methods using ground truth camera parameters as input. Particularly in scene-level scenarios, our method exhibits superior robustness. The error is reduced by at least 28.39% and 34.37% on DTU and Replica datasets respectively. Even without per-scene optimization, the surface reconstruction error is reduced by at least 14.62% and 21.93% respectively.
[…]
Scan 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
3DGS-GT 3.013 4.935 5.461 1.398 3.743 3.033 1.892 4.854 3.360 1.899 2.469 3.431 1.068 1.780 1.850 2.946
3DGS*-GT 2.010 4.382 5.277 1.873 5.771 3.885 2.035 4.676 4.197 1.819 5.241 3.914 1.507 5.866 3.613 3.738
2DGS-GT 5.654 7.908 7.011 7.401 8.661 5.950 4.638 9.286 5.231 5.641 7.449 5.809 3.290 4.765 6.797 6.366
2DGS*-GT 2.106 3.295 2.516 1.343 3.345 2.511 2.006 2.307 3.038 2.329 2.613 4.177 0.903 2.615 1.822 2.462
PGSR-GT 3.087 3.577 3.001 1.032 3.390 2.482 2.294 2.925 3.815 1.704 2.419 3.410 1.057 1.928 1.968 2.539
PGSR*-GT 3.360 3.280 2.750 1.270 5.150 1.840 0.880 1.790 3.490 1.190 1.860 1.110 0.610 1.090 1.520 2.079
FatesGS-GT 1.058 3.615 5.159 0.888 1.509 2.025 0.890 3.404 1.578 1.235 1.389 0.814 0.498 0.909 0.966 1.729
FatesGS*-GT 0.670 1.940 1.170 0.770 1.280 1.230 0.630 1.050 0.980 0.690 0.750 0.480 0.410 0.780 0.900 0.915
3DGS 6.999 9.218 8.181 6.916 10.86 6.262 5.188 10.26 5.483 5.447 7.216 6.438 3.969 5.242 6.725 6.961
3DGS* 5.759 9.947 8.383 7.127 9.431 6.293 5.480 8.755 6.106 6.507 8.211 7.454 4.245 6.184 7.421 7.154
2DGS 5.904 6.678 7.011 8.031 8.701 6.199 5.153 8.879 6.606 6.223 8.042 6.862 3.907 5.965 7.175 6.756
2DGS* 5.085 7.295 6.859 7.757 9.238 6.924 5.165 8.930 5.880 6.704 7.956 5.668 4.301 5.426 7.642 6.722
PGSR 5.359 6.676 6.628 7.168 7.922 5.913 4.733 8.571 5.532 5.646 7.188 4.797 3.037 4.569 6.534 6.018
PGSR* 3.745 8.030 6.021 7.004 8.612 6.006 4.574 9.085 4.618 6.163 7.511 3.879 3.226 4.687 6.718 5.992
FatesGS 2.506 7.510 4.145 2.222 4.233 4.949 3.179 4.417 6.048 2.532 3.806 3.507 2.901 2.594 2.805 3.824
[…]
CF-3DGS 8.488 9.307 8.346 9.740 8.800 8.132 6.818 8.248 8.919 7.755 8.442 7.437 7.841 7.860 8.009 8.276
FreeSplatter 5.673 8.894 9.206 5.903 7.602 9.477 4.660 4.587 8.468 3.274 4.352 7.873 3.436 3.867 4.461 6.116
[…]
Ours 1.300 2.103 1.767 1.468 2.066 1.476 1.339 2.191 1.736 1.479 1.419 1.420 0.793 1.572 1.583 1.581
Table 1: Quantitative experiments of CD on DTU datasets under different settings. We conduct experiments on all 15 scenes.
Scan Office0 Office1 Office2 Room0 Room1 Mean
3DGS-GT 183.52 179.71 172.41 177.93 176.67 178.05
3DGS*-GT 177.36 163.13 176.42 167.02 173.80 171.55
2DGS-GT 179.50 179.90 166.42 183.91 166.79 175.31
2DGS*-GT 152.44 144.69 187.64 181.86 168.51 167.03
PGSR-GT 168.11 183.27 167.57 181.68 174.74 175.08
PGSR*-GT 170.03 171.97 177.46 176.56 174.64 174.13
FatesGS-GT 181.22 176.32 183.77 178.73 188.13 181.64
FatesGS*-GT 187.27 153.78 183.85 186.56 175.29 177.35
3DGS 167.24 163.32 177.39 197.04 178.15 176.63
3DGS* 176.50 164.00 177.84 175.12 158.42 170.38
2DGS 181.87 149.37 196.53 176.93 165.06 173.95
2DGS* 176.00 162.06 172.93 174.21 161.86 169.41
PGSR 166.96 185.52 177.85 191.64 166.48 177.69
PGSR* 147.42 140.70 185.09 192.01 170.16 167.08
FatesGS 160.64 57.146 189.05 161.51 171.60 147.99
FatesGS* 110.32 95.371 183.05 181.56 153.97 144.85
CF-3DGS 172.16 165.08 152.55 173.52 184.27 169.52
FreeSplatter 169.90 158.22 172.16 163.02 173.96 167.45
[…]
Table 2: Quantitative experiments of CD on Replica datasets under different settings. We conduct experiments on 5 scenes.
Figure 5: Visualization of surface reconstruction on DTU and Replica datasets.
4.3 Evaluation of NVS
Subsequently, we evaluate the performance of our method on NVS. As demonstrated in Table 1 and Table 2, under conditions of sparse viewpoints and ground-truth camera parameters, simply increasing the number of iterations does not yield better reconstruction results; conversely, it leads to catastrophic scene geometry due to overfitting on the limited views. Therefore, we evaluate the NVS performance under equal iteration numbers. As shown in Table 3, the NVS performance of our method significantly surpasses other approaches. Even without per-scene optimization, where the scene is generated solely through end-to-end dense Gaussian generation, our method still demonstrates a clear advantage in NVS performance. Moreover, compared to the overfitting on limited views observed in methods represented by 3DGS, our approach effectively preserves scene geometry while performing per-scene optimization. The NVS error (LPIPS as an example) is reduced by at least 46.19% on DTU datasets and 73.13% on Replica datasets. Even without per-scene optimization, the error is decreased by at least 37.14% and 1.29% respectively.
3DGS
GT
2DGS
GT
PGSR
GT
FatesGS
GT
3DGS 2DGS PGSR FatesGS
CF
3DGS
Free
Splatter
VGGT
Ours
(wo Opt.)
Ours
[…]
Table 3: Quantitative evaluation of NVS on DTU and Replica datasets under different settings.
4.4 Ablation Study
To validate the effectiveness of each component in our proposed algorithm, we conducted comprehensive ablation studies. Specifically, we performed ablations on both the end-to-end dense Gaussian initialization and the geometrically-enhanced scene optimization, with the surface reconstruction quality evaluated using CD. Experiments were carried out on the DTU and Replica datasets. For per-scene optimization, we ablated the following components: monocular depth supervision $\mathcal{L}_{\text{rank}}$, multi-view feature consistency supervision $\mathcal{L}_{\text{MVS}}$, and differentiable camera poses $T^{\text{cam}}_{k}$. For Gaussian initialization, we ablated the contribution-based Gaussian pruning $Trim$ and the self-splitting-based Gaussian scene densification $\mathbb{D}$.
[…]
Replica DTU
Ours 23.87 1.581
No $\mathcal{L}_{\text{rank}}$ 24.66 2.905
No $\mathcal{L}_{\text{MVS}}$ 31.87 3.130
No $T^{\text{cam}}_{k}$ 32.29 2.855
Ours(wo Opt.) 27.28 2.208
No Trim 27.29 2.930
No $\mathbb{D}$ 44.21 3.584
Table 4: Ablation experiments on DTU and Replica datasets.
As shown in Table4, the removal of any individual component leads to a degradation in the surface reconstruction quality, thereby demonstrating the contribution of each module. It is worth noting that in the per-scene optimization stage, removing additional supervision may result in a reconstructed surface that is inferior to the initial scene. This can be primarily attributed to the geometric correctness of the dense Gaussian initialization and the inherent ambiguity in sparse viewpoints, which prevent simple increases in iteration counts from enhancing the scene geometry.
5 Conclusion
We propose FSFSplatter, a novel algorithm for rapid novel view synthesis and surface reconstruction from free-sparse RGB images. Our approach integrates Gaussian scene densification initialization and camera parameter estimation into an end-to-end large Transformer, which employs a self-splitting Gaussian head to map semi-dense point cloud features into a dense Gaussian scene, while applying contribution-based Gaussian pruning to eliminate useless floating artifacts. To address the issue of local view overfitting caused by sparse views, we incorporate supervision from monocular depth estimation and multi-view feature consistency to regularize the scene. Furthermore, we parameterize camera poses as differentiable variables to mitigate errors from camera parameter estimation. Compared to prior methods, FSFSplatter demonstrates superior accuracy and robustness, both with and without per-scene optimization. We achieve SOTA performance in free-sparse view surface reconstruction and NVS for both object-level and scene-level scenes, validated on the widely used DTU and Replica datasets.