### v2

# FSFSplatter: Build Surface and Novel Views with Sparse-Views within 2min

Abstract
Gaussian Splatting has become a leading reconstruction technique, known for its high-quality novel view synthesis and detailed reconstruction. However, most existing methods require dense, calibrated views. Reconstructing from free sparse images often leads to poor surface due to limited overlap and overfitting. We introduce FSFSplatter, a new approach for fast surface reconstruction from free sparse images. Our method integrates end-to-end dense Gaussian initialization, camera parameter estimation, and geometry-enhanced scene optimization. Specifically, FSFSplatter employs a large Transformer to encode multi-view images and generates a dense and geometrically consistent Gaussian scene initialization via a self-splitting Gaussian head. It eliminates local floaters through contribution-based pruning and mitigates overfitting to limited views by leveraging depth and multi-view feature supervision with differentiable camera parameters during rapid optimization. FSFSplatter outperforms current state-of-the-art methods on widely used DTU, Replica, and BlendedMVS datasets.
Figure 1: We address the inherent ambiguity and local view overfitting in the reconstruction of free-sparse images, proposing a novel pipeline named FSFSplatter that enables surface reconstruction and novel view synthesis within 2 minutes. FSFSplatter significantly reduces the number of iterations required while decreasing surface reconstruction error by at least 28.39% and novel view synthesis error by at least 46.19%. The figure illustrates the scene "scan63" from DTU datasets.
1 Introduction
Surface reconstruction and novel view synthesis (NVS) from unconstrained multi-view RGB images are long-standing tasks in autonomous driving[1], computer graphics[2], and virtual reality[3]. As typical methods require dense, calibrated images, accurate reconstruction and NVS from sparse, free inputs remain particularly challenging, yet crucial for scenarios with only limited RGB captures.
Traditional approaches often decompose the problem into separate stages of camera pose estimation, dense reconstruction, and surface extraction, typically by converting dense point clouds into surfaces via signed distance functions (SDFs). This pipeline, however, introduces errors at each stage, leading to irreversible error accumulation.
Recent studies have explored end-to-end models for unconstrained image reconstruction. For example, Dust3R[4] learns a robust image-to-PointMap mapping from image pairs, with several variants[5, 6]. Nevertheless, its network architecture is incapable of processing multiple views in a single forward pass, resulting in accumulated inconsistencies. VGGT[7] employs alternating global and local attention mechanisms, enabling 3D reconstruction from an arbitrary number of unconstrained images.
The sparsity of point clouds, however, makes them a suboptimal representation for many downstream applications (e.g., NVS). Recent breakthroughs in Neural Radiance Fields (NeRF)[8] and differentiable rendering[9] have demonstrated unprecedented quality in multi-view reconstruction. However, most of these methods struggle to generalize to unconstrained sparse-view settings, as insufficient image overlap often leads to the failure of Structure-from-Motion (SfM), and the inherent ambiguity in dense reconstruction arises from view extrapolation and occlusion under sparse views. Methods like PF-LRM[10, 11] employ a Transformer to map RGB images into a NeRF representation. Nevertheless, the surface bias of NeRF[12] and the separate point cloud head for camera pose estimation limit the final reconstruction quality. In contrast, Gaussian Splatting (GS) scenes, due to their explicit scene representation and unbiased geometry, have shown advantages in surface reconstruction[13, 2]. FreeSplatter[14] constructs a GS scene in a feed-forward manner from a single, unconstrained RGB image, and subsequently estimates camera parameters and aligns multiple views using an optimizer.
We employ a large transformer as the backbone network, which regresses scale-consistent camera parameters, depth maps, and a Gaussian scene via independent prediction heads. To predict dense and geometrically accurate Gaussian primitives, we specifically design a geometry and global feature-based Gaussian self-splitting decoder. A semi-dense Gaussian scene is first obtained by fusing the pixel-aligned depth map and DPT-decoded features. Subsequently, an Encoder-Decoder architecture is utilized to compute offset maps and perform Gaussian self-splitting, yielding the dense Gaussian scene. To enhance the overall geometric stability and eliminate local floaters, we apply a contribution-based pruning strategy to obtain the initial scene. Furthermore, to address the issue of local view overfitting during the optimization process under sparse inputs, we propose an optimization scheme supervised by monocular depth and multi-view features, which significantly improves both surface reconstruction and NVS quality within a short time.
Our main contributions can be summarized as follows:
  1. 1.
    We propose FSFSplatter, a fast surface reconstruction algorithm from free-sparse images. Compared to prior works, FSFSplatter achieves accurate reconstruction results both with and without per-scene optimization.
  2. 2.
    The Gaussian head performs Gaussian self-splitting based on high-dimensional features, coupled with contribution-based pruning, enabling precise initialization of the dense Gaussian scene. Furthermore, we introduce a geometry-enhanced scene optimization supervised by monocular depth and multi-view features, which effectively addresses the common issue of local view overfitting in sparse-view optimization.
  3. 3.
    We demonstrate state-of-the-art (SOTA) performance on widely-used DTU, Replica, and BlendedMVS datasets upon surface reconstruction and NVS.
2 Related Works
2.1 Surface Reconstruction
Surface reconstruction from RGB images is a fundamental problem in computer vision. Traditional methods typically represent scenes as point clouds [15], depth maps [16], or voxels [17]. In recent years, with the advancement of computational hardware and neural network technologies, implicit scene representations have been widely adopted. NeRF and volume rendering [8] have been applied to NVS and 3D reconstruction, undergoing rapid development. To address the surface bias inherent in volume rendering, NeuS [12, 18] proposed an unbiased NVS-based surface reconstruction method, achieving accurate surface reconstruction based on an implicit signed distance function (SDF) and demonstrating the advantages of NVS-based surface reconstruction in detailed regions.
3D Gaussian Splatting (3DGS) represents a scene using a set of explicit Gaussian primitives [9], enabling fast and high-quality NVS and reconstruction. However, 3DGS lacks expressive capacity for scene geometry. Subsequent work has achieved accurate surface reconstruction by constraining the shapes of Gaussian primitives to align with the surface [2, 13]. Through depth map rendering and fusion, fast and accurate geometric surface reconstruction can be realized. Furthermore, mapping algorithms from colored point clouds to Gaussians have further enhanced the reconstruction quality [19, 20].
2.2 Pose-free Reconstruciton
However, most of the aforementioned methods require dense camera views with known parameters. Under sparse views, Structure-from-Motion (SfM) often fails due to insufficient image overlap. Even with constrained camera parameters, errors in the camera parameters and sparsity can gradually lead to overfitting to individual views, resulting in erroneous geometric surfaces.
In traditional pipelines, image-based 3D reconstruction involves separate stages such as feature extraction, feature matching, bundle adjustment (BA), and plane sweeping. Colmap [15, 21], introduced in 2016, has become a benchmark method for sparse and dense reconstruction due to its exceptional robustness and performance.
However, both camera pose estimation from sparse views and NVS-based dense reconstruction methods suffer from unavoidable ambiguities caused by extrapolation and occlusion [22]. This leads the scene to overfit the sparse color observations, forming incorrect geometry that severely degrades the quality of surface reconstruction. Some approaches address this by leveraging large-scale dataset pre-training for better generalization to other datasets, but often require several days of training [22]. Methods focusing on per-scene optimization introduce additional supervision to enhance geometric stability during the optimization process [23, 24]. These methods typically incorporate monocular depth estimation [25] as a geometric prior to stabilize the optimization.
With the growing recognition of the potential of large Transformer models, pose estimation and 3D reconstruction are increasingly being implemented in an end-to-end manner, thereby mitigating the issue of cumulative errors inherent in the aforementioned multi-stage pipelines. Dust3R [4] and its subsequent developments [5, 6] achieve accurate camera parameter estimation and PointMap prediction from just two images. VGGT facilitates joint camera parameter estimation and 3D reconstruction from multiple images in a single forward pass by leveraging an alternating mechanism of global and local attention [7].
Beyond scene representations based on point clouds and depth maps, several works build upon NeRF and Gaussian Splatting (GS) representations. PF-LRM [10] maps unconstrained multi-view RGB images to a NeRF representation via a transformer, employing separate branches to predict point clouds and camera poses, which introduces additional training costs and challenges for surface reconstruction. FreeSplatter [14] maps unconstrained single views to a Gaussian scene representation and performs multi-view camera pose estimation and scene alignment based on optimization, treating scene generation and camera pose estimation as independent steps.
3 Method
We propose a fast reconstruction method based on free-sparse RGB images that enables robust 3D surface reconstruction from only three input images within 2 minutes.
Our framework consists of two main components: an end-to-end differentiable dense Gaussian initialization and camera parameter estimation, and a geometry-aware robust Gaussian scene optimization.
Figure 2: Overview of FSFSplatter. We employ a large alternating attention network as the backbone and generate high-dimensional semi-dense Gaussian scenes through independent heads. These are subsequently mapped into a dense Gaussian scene via a self-splitting densify MLP, forming an end-to-end framework for dense Gaussian initialization and camera parameter estimation. Through differentiable Gaussian scene construction from initialization and joint supervision of depth, multi-view stereo features, and RGBs during the optimization process, scene optimization achieves geometric enhancement, effectively mitigating error surfaces caused by overfitting to free-sparse views.
3.1 Preliminaries
3.1.1 Gaussian Splatting
Based on Gaussian Splatting [9], which represents a scene as a collection of Gaussian primitives, 2DGS further flattens these 3D Gaussians into 2D primitives [2], enabling precise surface reconstruction by closely adhering the Gaussians to the actual surfaces. Each Gaussian primitive is defined by its position $\mu^{3}$, rotation $R^{4}$, scale $S^{2}$, opacity $O^{1}$, and spherical harmonics (SH) coefficients $C^{48}$ for representing color.
The splatting of Gaussian primitives from the Gaussian coordinate system to the pixel coordinate system is achieved via a homogeneous transformation matrix $\mathbf{H}$, as shown in Equation 1.
$\mathbf{u}$ $=\mathbf{W}\mathbf{H}(s_{1},s_{2},1,1)^{\top}$ (1)
Multiple Gaussian primitives are composited via $\alpha$-blending, as shown in Equation 2. Here, $\mathcal{G}$ denotes the set of primitives influencing the pixel, and $\alpha$ is computed based on the Gaussian’s covariance matrix $\boldsymbol{\Sigma}$, its center coordinates $\boldsymbol{p}$, and the projected pixel coordinates $\boldsymbol{\mu}$.
$\mathbf{C}$ $=\sum_{i\in\mathcal{G}}\mathbf{c}_{i}\alpha_{i}\prod_{i-1}^{n}(1-\alpha_{n})$ (2)
$where\ \ \alpha_{n}$ $=\sigma_{n}\cdot\exp\left(-\frac{1}{2}\left(\boldsymbol{p}-\boldsymbol{\mu}_{n}\right)^{T}\boldsymbol{\Sigma}_{n}^{-1}\left(\boldsymbol{p}-\boldsymbol{\mu}_{n}\right)\right)$
3.1.2 Visual Geometry Grounded Transformer
To address the issue of error accumulation caused by multiple stages in traditional 3D reconstruction pipelines, VGGT [7] employs a large feed-forward neural network to directly infer all key attributes from multiple input views. These attributes include camera parameters $\mathbf{g}_{i}$, depth maps $D_{i}$, PointMaps $P_{i}$, and point trajectories $T_{i}$. The backbone network incorporates alternating global and local attention mechanisms. Utilizing a shared backbone with separate, dedicated prediction heads enables the simultaneous estimation of all required 3D attributes.
3.2 Problem definition
The input consists of three RGB images $I_{i}\in\mathbb{R}^{3\times 3\times H\times W}$ observing the same scene. Our algorithm $\mathcal{F}$ maps these images to a dense Gaussian representation of the scene, enabling accurate and fast novel view synthesis (NVS) and surface reconstruction through rasterization, as shown in Equation 3.
$\mathcal{F}(\left(I_{i}\right)^{3})=(\mathcal{G},(\mathcal{C}_{i})^{3})$ (3)
where $\mathcal{G}$ denotes the Gaussian scene. This work characterizes the Gaussian scene as a collection of 2D Gaussians within the 3D space, achieving precise alignment with surfaces and normals by flattening the 3D Gaussians [2]. Each Gaussian primitive comprises coordinates $X^{3}$, rotation $R^{3}$, scale $S^{2}$, opacity $O^{1}$, and spherical harmonic coefficients $\text{SH}^{48}$. $(\mathcal{C}_{i})^{3}$ represents the camera parameters (including both intrinsic and extrinsic parameters).
3.3 End-to-end dense Gaussian initialization and camera parameter estimation
Figure 3: The end-to-end dense Gaussian initialization and camera parameter estimation take sparse images as input. Image tokens are processed via a DINO Encoder and alternating attention mechanisms. A Camera Head and a DPT Head perform back-projection and depth estimation, yielding a semi-dense Gaussian scene. This scene is partitioned into multiple patches, which are then fed into a densification MLP to produce a dense Gaussian scene serving as the initialization.
Conventional 3D reconstruction pipelines are typically divided into stages such as feature extraction, feature matching, triangulation, and dense reconstruction. Each stage introduces errors that inevitably accumulate throughout the process.
Inspired by recent advances in 3D depth estimation [7, 4], we adopt a large Transformer architecture as the backbone for our end-to-end dense Gaussian initialization and camera parameter estimation framework. As illustrated in Figure 3, images are first encoded into tokens $(T_{i})^{3}$ using DINOv2 [26]. These tokens are then processed through the main network structure, which employs alternating global and local attention mechanisms.
Previous methods have demonstrated remarkable performance by incorporating decoder heads for camera parameters, depth estimation (DPT), and PointMap prediction [7]. To improve the initialization quality of the Gaussian Splatting (GS) scene, we propose a novel self-splitting-based patch initialization strategy for 2DGS scenes.
Figure 4: Explanation of dense Gaussian initialization. By incorporating explicit normal estimation and an MLP-based implicit Gaussian densification process, the quality of both novel NVS and surface reconstruction is significantly improved.
Based on the outputs from the camera head and the DPT head, we can rapidly perform back-projection to obtain pixel-aligned 3D point clouds. Concurrently, we design an independent DPT back-projection module that utilizes positional encoding to back-project the encoded tokens into pixel-aligned high-dimensional features $\mathcal{F}_{P}$. To address scale variations across different scenes, we first apply normalization to the point cloud. We employ an encoder-decoder architecture to map the aforementioned information into dense Gaussian primitive attributes. Specifically, PointMLP [27] is adopted to encode the geometric information of the point cloud, yielding high-dimensional geometric features $\mathcal{F}^{c}$. Subsequently, $\mathcal{F}_{P}$ is concatenated channel-wise with explicit Gaussian attributes—including $\text{SH}^{48}_{P}$, $R^{4}_{P}$, $S^{2}_{P}$, and $\mathcal{F}_{P}$—to form the encoded representation. We then design mutually independent decoders for the attribute variations $\Delta\mathcal{G}$ of all Gaussian primitives. These decoders predict the densification tendency of the original semi-dense Gaussian scene, as visualized in Fig. 4. Each Gaussian primitive splits into $N$ new primitives based on its predicted variation, as formulated in Equation 4.
$\Delta\mathcal{G}_{N\cdot P}=\{\Delta X^{3}_{N\cdot P},\Delta R^{4}_{N\cdot P},\Delta S^{2}_{N\cdot P},\Delta O^{1}_{N\cdot P},$ $\Delta\text{SH}^{48}_{N\cdot P}\}=D(\text{Cat}(\mathcal{F}_{P},\text{SH}^{48}_{P},R^{4}_{P},S^{2}_{P},\mathcal{F}_{P}))$ (4)
$\text{Densify}(\mathcal{G}_{P})$ $=\mathcal{G}_{P}+\Delta\mathcal{G}_{N\cdot P}$
Drawing on advanced point cloud processing techniques [28, 20], we adopt point cloud patches as the input units to our network to enhance its generalization capability across diverse scenes. After obtaining the back-projected 3D point cloud via the DPT head and camera head, and applying normalization, we employ K-Nearest Neighbors (KNN) to partition the point cloud into randomly sampled local patches. These patches are divided into multiple small batches, each processed through an encoder-decoder architecture to generate densified Gaussian primitives. Finally, after de-normalization, the processed batches are merged to form the complete Gaussian scene, as expressed in Equation 5. Here, $\mathbf{N}$ and $\mathbf{N}^{\prime}$ denote the normalization and de-normalization operations, respectively; $\mathbb{B}$ and $\mathbb{B}^{\prime}$ represent batch partitioning and merging; and $\mathbb{D}$ signifies the aforementioned densification process.
$\mathcal{G}=\mathbf{N}^{\prime}\{\mathbb{B}^{\prime}[\mathbb{B}(\mathbf{N}(\mathcal{G}_{init}))+\mathbb{D}\left(\mathbb{B}\left(\mathbf{N}\left(\mathcal{G}_{init}\right)\right)\right)]\}$ (5)
3.3.1 Training strategy
Our entire pipeline is fully end-to-end differentiable, enabling comprehensive training based on $\mathcal{L}_{\text{RGB-D}}$ and $\mathcal{L}_{\text{Cam}}$. Here, $\mathcal{L}_{\text{RGB-D}}$ comprises an RGB-based loss combining $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$, along with a depth-based loss $\mathcal{L}_{\text{D}}$. The camera parameters are defined by the intrinsic matrix $\mathbf{F}\in\mathbb{R}^{1}$ and the extrinsic parameters $\mathbf{T}=[\mathbf{t}\in\mathbb{R}^{3},\mathbf{q}\in\mathbb{R}^{4}]$, forming the camera loss $\mathcal{L}_{\text{Cam}}=\mathcal{L}_{\mathbf{F}}+\mathcal{L}_{\mathbf{t}}$.
A straightforward approach involves training all parameters of the entire pipeline directly. However, in practice, since both the input and the final pipeline output are RGB images, this often leads to overfitting on the novel view synthesis (NVS) of local views during training. This overfitting can result in erroneous geometry and degrade the quality of surface reconstruction [14]. To mitigate this, we adopt a geometry-progressive training strategy. We initialize our network backbone with pre-trained weights from VGGT-1B [7]. On this foundation, we remove the PointMap Head and Tracking Head, and proceed to train the DPT Head, Camera Head, and the backbone network. During this stage, the losses $\mathcal{L}_{\text{D}}$ and $\mathcal{L}_{\text{Cam}}$ are enabled to ensure geometric stability, maintaining consistency with the training methodology of VGGT.
To achieve a high-quality initialization of the Gaussian scene and prevent overfitting commonly associated with end-to-end NVS training, we subsequently freeze the DPT Head, Camera Head, and the backbone network. This freezing preserves geometric stability during the subsequent training phase. The DPT unproject Head and the densification MLP are then trained independently. In this stage, KNN Patch partitioning is disabled, allowing for global densification and estimation of Gaussian primitives across the entire scene, which serves as a form of pre-training. The losses $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ are enabled during this phase.
Finally, we unfreeze all parameters of the pipeline, enabling the backbone network to learn the mapping from RGB images to Gaussian attributes directly. The KNN Patch operation is reinstated during the densification process to enhance generalization. The complete set of loss functions is activated throughout this final training stage.
3.4 Geometrically-enhanced scene optimization
Figure 5: After initialization, a contribution-based Gaussian pruning process is first applied to eliminate erroneous floaters present in the dense Gaussians. Both the Gaussian primitives and the camera poses are treated as differentiable variables. Multi-view RGB, depth, and normal maps are rendered through rasterization. Various loss functions are constructed based on monocular depth estimation and multi-view high-dimensional feature extraction, with subsequent backpropagation.
3.4.1 Differentiable Gaussian scene construction
2D/3DGS jointly optimizes NVS and geometry through RGB-based losses [2, 9]. However, due to sparse viewpoint coverage and unknown camera parameters, direct optimization using RGB-based losses introduces inevitable multi-view ambiguity [22]. This often leads to color overfitting on sparse views and results in catastrophic geometry degradation. To address this, we propose a geometry-enhanced optimization strategy for Gaussian scenes.
The dense Gaussian scene initialization yields Gaussian primitives that are more densely distributed than pixels. While this provides a strong prior, it also introduces numerous occluded or invisible primitives. During optimization, these non-contributing primitives do not receive gradient updates through backpropagation, making them impossible to prune simply by filtering low-opacity elements. Such primitives are geometrically meaningless and often detrimental to surface reconstruction. Therefore, we perform contribution-based Gaussian pruning prior to optimization.
Referring to Equation 2, the contribution $\mathbf{C}_{n}$ of a Gaussian primitive $\mathcal{G}n$ can be quantified by its weight in the $\alpha$-blending process: $\alpha_{i}\prod{i-1}^{n}(1-\alpha_{n})$, as shown in Equation 6. Here, $\mathcal{P}_{n}$ denotes the set of pixels influenced by the Gaussian primitive $\mathcal{G}_{n}$. To prevent disproportionately large contributions from Gaussians covering extensive areas, we normalize the contribution by the number of occupied pixels $|\mathcal{P}_{n}|$ during accumulation. After computing contributions for all primitives, we sort them and prune those with low contribution or excessively low opacity, as shown in Figure 6.
$\mathbf{C}_{n}=\sum_{k=1}^{3}\frac{1}{|\mathcal{P}_{n}|}\sum_{p\in\mathcal{P}_{n}}(\alpha_{i}^{p}\prod_{i-1}^{n}(1-\alpha_{n}^{p}))$ (6)
Figure 6: Contribution-based Gaussian pruning for scene initialization
2DGS employs a CUDA-based differentiable rasterizer that explicitly computes the gradients of Gaussian primitives. However, the original rasterizer does not support the backward propagation of camera parameters, which is detrimental for pose-free 3D reconstruction, as it tends to amplify subtle multi-view inconsistencies during the optimization process. To address this limitation, we incorporate differentiable camera poses into the differentiable Gaussian primitives. Following related works such as GS-SLAM [29], we consistently position the camera pose at the origin of the coordinate system. Prior to rasterization, we apply the inverse transformation of the camera pose $T^{\text{cam}}_{k}$ to all Gaussian primitives, thereby achieving the same view synthesis effect as in the original framework. Throughout this process, $T^{\text{cam}}_{k}$ remains differentiable. Consequently, any loss based on the rasterized image will propagate back to the camera pose before being passed to the Gaussian primitives during backpropagation. Owing to the sparsity of viewpoints, our approach enables rapid and mutually independent pose optimization across different views.
3.4.2 Loss Function
During the optimization process, our loss function comprises several components: the RGB-based $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$, the normal map-based $\mathcal{L}_{n}$, the losses derived from the rendered depth map and monocular depth estimation-namely $\mathcal{L}_{\text{rank}}$, $\mathcal{L}_{\text{smooth}}$, and $\mathcal{L}_{d}$-as well as $\mathcal{L}_{\text{MVS}}$, which is based on RGB feature extraction and rendered depth back-projection, as illustrated in Figure5. Among these, $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{SSIM}}$ are consistent with 3DGS [9]; $\mathcal{L}_{d}$ and $\mathcal{L}_{n}$ are are consistent with 2DGS [2]. Furthermore, owing to the dense initialization of the Gaussian scene, we are able to apply all loss supervision from the beginning, without the need for a gradual progression starting solely from the RGB Loss.
We employ Marigold, a Vision Transformer-based monocular depth estimation network, as the geometric prior [30]. The generalized monocular depth estimation $D_{\text{est}}$ ensures the overall stability of the scene during optimization. The 2D Gaussian primitives achieve accurate depth map rendering by closely adhering to the actual surfaces. To address the scale ambiguity inherent in monocular depth estimation, we utilize a depth ranking loss to supervise the rasterized depth $D_{\text{re}}$, as shown in Equation 7. Specifically, pixel coordinates $p_{1},p_{2}$ are sampled from neighboring patches $\mathcal{P}$ using the same random strategy $\mathcal{R}$ applied to both $D_{\text{est}}$ and $D_{\text{re}}$, in order to compute the depth difference. Here, sgn denotes the sign function, and $\sigma$ represents the ReLU function.
$\mathcal{L}_{\text{rank}}=\sum_{p_{1},p_{2}\in\mathcal{P}}\sigma\left(\operatorname{sgn}(\mathcal{S}(D_{\text{pre}},p_{1},p_{2})\cdot\mathcal{S}(D_{\text{re}},p_{1},p_{2}))+m\right)$ (7)
$where\ \ \ \mathcal{S}(D,p_{1},p_{2})=D(p_{1})-D(p_{2})$
Concurrently, we enforce a depth smoothness loss $\mathcal{L}_{\text{smooth}}$ in regions where the monocular depth estimation does not exhibit rapid variations, to preserve the smoothness of the synthesized surfaces. As shown in Equation 8, where $n_{1}$ and $n_{2}$ are small thresholds used to identify edges in the monocular depth estimate.
$\mathcal{L}_{\text{smooth}}=\sum_{p_{1},p_{2}\in\mathcal{P}}\text{sgn}(|D_{\text{pre}}(p_{1})-D_{\text{pre}}(p_{2})|-n_{1})\cdot\sigma((|D_{\text{re}}(p_{1})-D_{\text{re}}(p_{2})|-n_{2})$ (8)
Loss based on monocular depth estimation effectively preserves the overall geometric structure of the scene, yet often lacks sufficient supervision for fine details. While direct RGB supervision of novel views could theoretically provide detailed guidance, inconsistencies in multi-view illumination and camera capture variations introduce color discrepancies across different views [31]. These discrepancies are further exacerbated in free-sparse views, leading to inconsistencies during training.
To address this issue, we employ a multi-view feature alignment loss based on a U-Net architecture, which supervises high-dimensional features derived from both RGB data and rendered depth back-projection. Specifically, we extract high-dimensional features from the input RGB images using a pre-trained U-Net [32]. These features, along with the rendered depth map, are back-projected to reference views. The feature difference between the current view and reference views is then computed as shown in Equation 9.
$\mathcal{L}_{\text{MVS}}=\sum_{r}\left|1-\cos\left[\mathcal{R_{r}}\left(\mathbf{U}^{(c)}\left(I_{r}\right)\right),\mathbf{U}^{(c)}\left(I\right)\right]\right|$ (9)
During training, to accelerate the optimization process, we control the downsampling resolution based on image frequency [33]. Unlike sparse point clouds initialized by Colmap, the dense Gaussian initialization allows the initial training downsampling rate to be set to 2. Owing to the sub-pixel level density of the initial Gaussians, we do not perform gradient-based densification during optimization. Instead, we only execute Gaussian pruning based on opacity and contribution to eliminate low contributing Gaussian primitives and error floaters.
4 Experiments
4.1 Settings
Training
Our Gaussian initialization parameter $K$ is set to 2, with an image resolution of $512\times 512$ and a single patch size of 2048. All our experiments are conducted on a server equipped with an Intel Platinum 8470Q processor and an Nvidia RTX5090 GPU, requiring 9 hours to complete 120 training epochs. The training dataset consists of BlendMVS [34], DTU [35], Virtual KITTI [36], and Replica [37], encompassing extensive RGB-D data and camera parameters from both synthetic and real-captured sources.
Metrics
We employ widely adopted metrics including PSNR, SSIM[38], and LPIPS[39] to evaluate NVS quality; employ Chamfer Distance (CD)[35] to evaluate surface reconstruction quality; employ ATE RMSE[40] (rotation and translation) to evaluate camera pose estimation.
Baselines
We evaluate the reconstruction results on base methods: 3D Gaussian Splatting 3DGS [9]; surface reconstruction method 2DGS [2]. Pose-free methods: RGB based pose-free reconstruction method CF-3DGS [41]; free-sparse-view feed-forward Gaussian Splatting generation method FreeSplatter [14]; dense reconstruction method based on large Transformer VGGT. Sparse-view surface reconstruction methods: sparse-view surface reconstruction method with additional regularization FatesGS; surface reconstruction method incorporating unbiased depth estimation PSGR.
Among these algorithms, 3DGS, 2DGS, FatesGS, and PSGR, as classical GS-based per-scene optimization pipelines, require camera parameters as prior knowledge. We therefore employ VGGT to estimate the camera parameters and subsequently evaluate the performance of these methods under both ground truth camera parameters and estimated camera parameters. Furthermore, we conduct comparative evaluations under two training configurations: one using our fixed iteration count of 1000, and the other using their default iteration settings.
Datasets
We evaluate FSFSplatter on DTU datasets [35], Replica datasets [37] and BlendedMVS datasets[34]. DTU datasets have been extensively used in prior surface reconstruction works. We evaluate on all 15 scenes from DTU, where each scene contains 49 or 64 images at a resolution of $1600\times 1200$. Following previous works [23, 42], we select views 23, 24, and 33 as input and downsample the images to $800\times 600$. Since DTU datasets primarily consist of object-level scenes, we further validate our algorithm on Replica and BlendedMVS datasets to assess its generalization capability and performance on scene-level scenes, which respectively contain simulated images and real-captured images. We respectively select the five scenes and randomly choose three input views at a resolution of $1200\times 680$ and $768\times 576$.
Experimental Platform
Experiments were carried out utilizing an NVIDIA RTX 5090 GPU alongside an Intel Platinum 8470Q CPU. The software setup operated on Ubuntu 20.04, with PyTorch version 2.3.1 and CUDA version 12.1.
4.2 Evaluation of Surface Reconstruction
We first evaluate the surface reconstruction accuracy (CD) under different settings across all baselines. As shown in Table1, Table2, and Figure7, FSFSplatter demonstrates significant advantages in CD error even when compared with per-scene optimization methods using their default iteration numbers, and remains competitive with methods using ground truth camera parameters as input. Particularly in scene-level scenarios, FSFSplatter exhibits superior robustness. The error is reduced by at least 28.39% and 34.37% on DTU and Replica datasets respectively. Even without per-scene optimization, the surface reconstruction error is reduced by at least 14.62% and 21.93% respectively.
[…]
Scan 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
3DGS-GT 3.013 4.935 5.461 1.398 3.743 3.033 1.892 4.854 3.360 1.899 2.469 3.431 1.068 1.780 1.850 2.946
3DGS*-GT 2.010 4.382 5.277 1.873 5.771 3.885 2.035 4.676 4.197 1.819 5.241 3.914 1.507 5.866 3.613 3.738
2DGS-GT 5.654 7.908 7.011 7.401 8.661 5.950 4.638 9.286 5.231 5.641 7.449 5.809 3.290 4.765 6.797 6.366
2DGS*-GT 2.106 3.295 2.516 1.343 3.345 2.511 2.006 2.307 3.038 2.329 2.613 4.177 0.903 2.615 1.822 2.462
PGSR-GT 3.087 3.577 3.001 1.032 3.390 2.482 2.294 2.925 3.815 1.704 2.419 3.410 1.057 1.928 1.968 2.539
PGSR*-GT 3.360 3.280 2.750 1.270 5.150 1.840 0.880 1.790 3.490 1.190 1.860 1.110 0.610 1.090 1.520 2.079
FatesGS-GT 1.058 3.615 5.159 0.888 1.509 2.025 0.890 3.404 1.578 1.235 1.389 0.814 0.498 0.909 0.966 1.729
FatesGS*-GT 0.670 1.940 1.170 0.770 1.280 1.230 0.630 1.050 0.980 0.690 0.750 0.480 0.410 0.780 0.900 0.915
3DGS 6.999 9.218 8.181 6.916 10.86 6.262 5.188 10.26 5.483 5.447 7.216 6.438 3.969 5.242 6.725 6.961
3DGS* 5.759 9.947 8.383 7.127 9.431 6.293 5.480 8.755 6.106 6.507 8.211 7.454 4.245 6.184 7.421 7.154
2DGS 5.904 6.678 7.011 8.031 8.701 6.199 5.153 8.879 6.606 6.223 8.042 6.862 3.907 5.965 7.175 6.756
2DGS* 5.085 7.295 6.859 7.757 9.238 6.924 5.165 8.930 5.880 6.704 7.956 5.668 4.301 5.426 7.642 6.722
PGSR 5.359 6.676 6.628 7.168 7.922 5.913 4.733 8.571 5.532 5.646 7.188 4.797 3.037 4.569 6.534 6.018
PGSR* 3.745 8.030 6.021 7.004 8.612 6.006 4.574 9.085 4.618 6.163 7.511 3.879 3.226 4.687 6.718 5.992
FatesGS 2.506 7.510 4.145 2.222 4.233 4.949 3.179 4.417 6.048 2.532 3.806 3.507 2.901 2.594 2.805 3.824
FatesGS* 3.039 7.317 6.952 1.915 5.534 5.071 4.304 4.237 4.332 2.370 3.041 2.607 2.220 1.981 2.919 3.856
CF-3DGS 8.488 9.307 8.346 9.740 8.800 8.132 6.818 8.248 8.919 7.755 8.442 7.437 7.841 7.860 8.009 8.276
FreeSplatter 5.673 8.894 9.206 5.903 7.602 9.477 4.660 4.587 8.468 3.274 4.352 7.873 3.436 3.867 4.461 6.116
[…]
Ours 1.300 2.103 1.767 1.468 2.066 1.476 1.339 2.191 1.736 1.479 1.419 1.420 0.793 1.572 1.583 1.581
Table 1: Quantitative experiments of CD ($m$) on DTU datasets under different settings. We conduct experiments on all 15 scenes.
Scan Office0 Office1 Office2 Room0 Room1 Mean
3DGS-GT 183.52 179.71 172.41 177.93 176.67 178.05
3DGS*-GT 177.36 163.13 176.42 167.02 173.80 171.55
2DGS-GT 179.50 179.90 166.42 183.91 166.79 175.31
2DGS*-GT 152.44 144.69 187.64 181.86 168.51 167.03
PGSR-GT 168.11 183.27 167.57 181.68 174.74 175.08
PGSR*-GT 170.03 171.97 177.46 176.56 174.64 174.13
FatesGS-GT 181.22 176.32 183.77 178.73 188.13 181.64
FatesGS*-GT 187.27 153.78 183.85 186.56 175.29 177.35
3DGS 167.24 163.32 177.39 197.04 178.15 176.63
3DGS* 176.50 164.00 177.84 175.12 158.42 170.38
2DGS 181.87 149.37 196.53 176.93 165.06 173.95
2DGS* 176.00 162.06 172.93 174.21 161.86 169.41
PGSR 166.96 185.52 177.85 191.64 166.48 177.69
PGSR* 147.42 140.70 185.09 192.01 170.16 167.08
FatesGS 160.64 57.146 189.05 161.51 171.60 147.99
FatesGS* 110.32 95.371 183.05 181.56 153.97 144.85
CF-3DGS 172.16 165.08 152.55 173.52 184.27 169.52
FreeSplatter 169.90 158.22 172.16 163.02 173.96 167.45
[…]
Table 2: Quantitative experiments of CD ($m$) on Replica datasets under different settings. We conduct experiments on 5 scenes.
Figure 7: Visualization of surface reconstruction on DTU and Replica datasets.
4.3 Evaluation of NVS
Subsequently, we evaluate the performance of FSFSplatter on NVS. As demonstrated in Table 1 and Table 2, under conditions of sparse viewpoints and ground-truth camera parameters, simply increasing the number of iterations does not yield better reconstruction results; conversely, it leads to catastrophic scene geometry due to overfitting on the limited views. Therefore, we evaluate the NVS performance under equal iteration numbers. As shown in Table 3 and Figure 8, the NVS performance of FSFSplatter significantly surpasses other approaches. Even without per-scene optimization, where the scene is generated solely through end-to-end dense Gaussian generation, FSFSplatter still demonstrates a clear advantage in NVS performance. Moreover, compared to the overfitting on limited views observed in methods represented by 3DGS, our approach effectively preserves scene geometry while performing per-scene optimization. The NVS error (LPIPS as an example) is reduced by at least 46.19% on DTU datasets, 73.13% on Replica datasets, and 87.35% on BlendedMVS datasets. Even without per-scene optimization, the error is decreased by at least 37.14%, 1.29%, and 14.36% respectively.
Figure 8: Visualization of NVS on DTU, Replica, and BlendedMVS datasets.
3DGS
GT
2DGS
GT
PGSR
GT
FatesGS
GT
3DGS 2DGS PGSR FatesGS
CF
3DGS
Free
Splatter
VGGT
Ours
(wo Opt.)
Ours
[…]
Table 3: Quantitative evaluation of NVS on DTU, Replica, and BlendedMVS datasets under different settings.
4.4 Evaluation of pose extimation
As a surface reconstruction and NVS method based on free-sparse images, we also evaluated the accuracy of camera pose estimation on the Replica and DTU datasets. Since both FSFSplatter and the pipelines are based on RGB images, we aligned the scale by registering the camera poses following established works [4].
As shown in Table 4, FSFSplatter achieves competitive performance in terms of ATE RMSE compared to SOTA pose estimation methods, while simultaneously delivering high-quality surface reconstruction and NVS results.
CF-3DGS FreeSplatter Fast3R Dust3r VGGT Ours
[…]
Table 4: Pose estimation evaluation ATE RMSE on Replica and DTU datasets.
4.5 Evaluation of optimization speed
As a reconstruction method that supports per-scene optimization, FSFSplatter achieves high reconstruction accuracy from the initial stage due to its well-structured Gaussian dense initialization. Consequently, FSFSplatter attains remarkably short optimization iteration times. As shown in Table 5, we evaluate the optimization speed of FSFSplatter and other per-scene optimization-based reconstruction algorithms on Replica datasets.
Method Time(s)
3DGS 429.45
2DGS 840.37
FatesGS 915.93
PGSR 1018.2
Ours 107.39
Table 5: Optimization speed evaluation of per-scene optimization-based algorithms on Replica datasets.
4.6 Ablation study
To validate the effectiveness of each component in our proposed algorithm, we conducted comprehensive ablation studies. Specifically, we performed ablations on both the end-to-end dense Gaussian initialization and the geometrically-enhanced scene optimization, with the surface reconstruction quality evaluated using CD. Experiments were carried out on the DTU and Replica datasets. For per-scene optimization, we ablated the following components: monocular depth supervision $\mathcal{L}_{\text{rank}}$, depth smoothness loss $\mathcal{L}_{\text{smooth}}$, multi-view feature consistency supervision $\mathcal{L}_{\text{MVS}}$, and differentiable camera poses $T^{\text{cam}}_{k}$. For Gaussian initialization, we ablated the contribution-based Gaussian pruning $Trim$ and the self-splitting-based Gaussian scene densification $\mathbb{D}$.
[…]
Replica DTU
Ours 23.87 1.581
No $\mathcal{L}_{\text{rank}}$ 24.66 2.905
No $\mathcal{L}_{\text{smooth}}$ 37.71 1.895
No $\mathcal{L}_{\text{MVS}}$ 31.87 3.130
No $T^{\text{cam}}_{k}$ 32.29 2.855
Ours(wo Opt.) 27.28 2.208
No Trim 27.29 2.930
No $\mathbb{D}$ 44.21 3.584
Table 6: Ablation experiments on DTU and Replica datasets.
As shown in Table6, the removal of any individual component leads to a degradation in the surface reconstruction quality, thereby demonstrating the contribution of each module. It is worth noting that in the per-scene optimization stage, removing additional supervision may result in a reconstructed surface that is inferior to the initial scene. This can be primarily attributed to the geometric correctness of the dense Gaussian initialization and the inherent ambiguity in sparse viewpoints, which prevent simple increases in iteration counts from enhancing the scene geometry.
5 Conclusion
We propose FSFSplatter, a novel algorithm for rapid novel view synthesis and surface reconstruction from free-sparse RGB images. Our approach integrates Gaussian scene densification initialization and camera parameter estimation into an end-to-end large Transformer, which employs a self-splitting Gaussian head to map semi-dense point cloud features into a dense Gaussian scene, while applying contribution-based Gaussian pruning to eliminate useless floating artifacts. To address the issue of local view overfitting caused by sparse views, we incorporate supervision from monocular depth estimation and multi-view feature consistency to regularize the scene. Furthermore, we parameterize camera poses as differentiable variables to mitigate errors from camera parameter estimation. Compared to prior methods, FSFSplatter demonstrates superior accuracy and robustness, both with and without per-scene optimization. We achieve SOTA performance in free-sparse view surface reconstruction and NVS for both object-level and scene-level scenes, validated on the widely used DTU, Replica, and BlendedMVS datasets.
Appendix A Details of framework and implementation
A.1 Implementation
The scene representation of FSFSplatter is based on 2DGS, and the rasterization pipeline is implemented based on AbsGS [43] and TrimGS [44]. Compared to basic 2DGS pipeline [2], the rasterization pipeline of FSFSplatter employs Abs accumulated gradient $\bar{\Delta}_{i}^{\text{ACC}}$ when caculating the accumulated gradient ${\Delta}_{i}^{\text{ACC}}$ (i.e., the gradient of the loss function $\mathcal{L}$ with respect to the projected pixel coordinate $u$). This approach mitigates gradient collision, thereby enabling accurate identification of suboptimal Gaussian primitives and addressing the blurring and loss of detail caused by over-reconstruction during Gaussian splitting, as shown in Equation S1. Here, $m$ denotes the sequence of all affected pixels. Contribution-based pruning is implemented following TrimGS, with the key distinction being the extension from pruning designed for NVS in 3DGS to pruning designed for surface reconstruction in 2DGS, as formulated in Equation 6.
$\bar{\Delta}_{i}^{\text{ACC}}=\sum_{j=1}^{m}\left|\frac{\partial\mathcal{L}_{j}}{\partial u_{i}}\right|,\quad{\Delta}_{i}^{\text{ACC}}=\sum_{j=1}^{m}\frac{\partial\mathcal{L}_{j}}{\partial u_{i}}$ (S1)
A.2 Parameter Setting
For the contribution-based Gaussian pruning, we perform a pruning at the start of optimization and subsequently repeat it every 500 iterations, removing the 5% of Gaussian primitives with the lowest contribution. For the loss function, the parameters $m$, $n_{1}$, and $n_{2}$ are set to $\times 10^{-4}, 1×1$$\times 10^{-2}$, and $1\times 10^{-4}$, respectively.
A.3 Evaluation Metrics
We employ PSNR, SSIM, and LPIPS to comprehensively evaluate the quality of NVS. For DTU datasets [35] and BlendedMVS datasets [34], where transparent masks are provided, we apply masks to all images prior to evaluation to ensure fair comparisons. To assess surface reconstruction quality, we utilize the Chamfer Distance (CD). Following previous works [35, 23], we adopt the L2 Chamfer Distance as our evaluation metric, as shown in Equation S2.
For evaluating camera pose estimation quality, we use the ATE RMSE. Since FSFSplatter is entirely based on RGB images, estimated poses differ from the ground truth by a scale factor. Following prior research [40], we perform rigid registration on the camera poses before evaluation and compute the absolute error in the ground truth scale. This approach eliminates inconsistencies in error evaluation caused by scale differences.
$\mathrm{CD}\left(\mathcal{M},\mathcal{M}_{\text{gt}}\right)=\frac{1}{2\left|\mathcal{M}\right|}\sum_{{x}\in\mathcal{M}}\min_{{y}\in\mathcal{M}_{\text{gt}}}\|{x}-{y}\|_{2}+\frac{1}{2\left|\mathcal{M}_{\text{gt}}\right|}\sum_{{y}\in\mathcal{M}_{\text{gt}}}\min_{{x}\in\mathcal{M}}\|{x}-{y}\|_{2}$ (S2)
Appendix B Detailed experiments
B.1 Detailed NVS quantization experiments
Table 3 presents the average quantitative evaluation results of NVS on DTU, Replica, and BlendedMVS datasets, while Tables S1, Table S2, and Table S3 provide the detailed per-scene evaluation results for each dataset.
Scan 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean
PSNR 31.128 29.295 25.873 36.033 34.249 32.639 32.441 37.970 31.912 36.649 37.903 34.387 34.111 40.040 40.704 34.356
3DGS-GT SSIM 0.841 0.887 0.865 0.970 0.949 0.964 0.935 0.963 0.937 0.933 0.951 0.945 0.931 0.959 0.969 0.933
LPIPS 0.279 0.158 0.231 0.080 0.118 0.094 0.199 0.103 0.152 0.157 0.158 0.164 0.173 0.132 0.098 0.153
PSNR 20.265 22.157 23.860 28.613 26.711 25.606 26.693 30.251 25.402 28.982 31.817 27.723 30.078 36.417 35.506 28.006
2DGS-GT SSIM 0.609 0.782 0.745 0.907 0.927 0.937 0.882 0.941 0.869 0.895 0.923 0.903 0.886 0.940 0.953 0.873
LPIPS 0.489 0.230 0.326 0.154 0.144 0.134 0.276 0.130 0.232 0.204 0.193 0.216 0.226 0.158 0.118 0.215
PSNR 27.518 26.766 25.185 35.262 28.352 27.879 29.146 30.967 27.968 30.423 30.004 32.456 34.668 40.089 38.549 31.016
PGSR-GT SSIM 0.841 0.905 0.906 0.978 0.944 0.963 0.911 0.950 0.922 0.937 0.954 0.937 0.957 0.973 0.976 0.937
LPIPS 0.251 0.122 0.156 0.046 0.089 0.071 0.149 0.096 0.115 0.119 0.113 0.143 0.081 0.064 0.054 0.111
PSNR 25.790 17.563 14.710 26.701 14.782 24.567 23.939 17.852 20.589 17.613 27.563 25.467 27.786 30.111 29.991 23.002
FatesGS-GT SSIM 0.780 0.703 0.800 0.795 0.804 0.816 0.804 0.725 0.800 0.801 0.849 0.792 0.844 0.861 0.842 0.801
LPIPS 0.343 0.281 0.280 0.315 0.240 0.374 0.333 0.438 0.323 0.375 0.351 0.400 0.296 0.332 0.358 0.336
PSNR 20.938 23.568 25.674 28.973 29.863 28.280 27.256 34.471 26.401 32.361 30.170 28.248 30.539 36.792 37.067 29.373
3DGS SSIM 0.618 0.787 0.730 0.906 0.933 0.943 0.894 0.954 0.878 0.901 0.928 0.909 0.889 0.941 0.955 0.878
LPIPS 0.487 0.226 0.329 0.149 0.140 0.128 0.261 0.118 0.225 0.201 0.186 0.201 0.225 0.158 0.118 0.210
PSNR 20.212 16.725 21.856 28.686 26.678 25.487 26.283 27.593 22.665 28.945 30.618 26.986 30.025 30.796 35.225 26.585
2DGS SSIM 0.603 0.724 0.693 0.907 0.926 0.932 0.876 0.929 0.820 0.894 0.924 0.900 0.885 0.910 0.950 0.858
LPIPS 0.495 0.267 0.355 0.154 0.145 0.140 0.290 0.143 0.292 0.205 0.192 0.219 0.228 0.197 0.124 0.230
PSNR 21.146 17.712 21.858 29.532 26.743 24.684 26.983 27.803 22.494 28.793 30.880 28.184 31.726 32.199 36.279 27.134
PGSR SSIM 0.636 0.716 0.733 0.941 0.929 0.934 0.876 0.926 0.811 0.912 0.944 0.908 0.928 0.922 0.965 0.872
LPIPS 0.463 0.289 0.326 0.094 0.106 0.112 0.195 0.138 0.256 0.158 0.125 0.184 0.121 0.144 0.069 0.185
PSNR 16.564 16.781 20.542 23.955 29.153 20.238 23.836 24.370 20.296 22.916 25.261 24.491 26.948 31.785 29.240 23.758
FatesGS SSIM 0.506 0.729 0.662 0.828 0.947 0.889 0.798 0.894 0.783 0.833 0.876 0.867 0.844 0.911 0.916 0.819
LPIPS 0.592 0.273 0.376 0.210 0.090 0.159 0.323 0.170 0.287 0.244 0.231 0.244 0.238 0.168 0.141 0.250
PSNR 19.495 17.156 21.760 24.198 24.421 27.438 27.590 29.752 23.873 26.474 28.559 28.852 27.106 30.780 32.207 25.977
CF-3DGS SSIM 0.650 0.754 0.715 0.839 0.918 0.933 0.868 0.949 0.845 0.883 0.884 0.917 0.859 0.904 0.927 0.856
LPIPS 0.328 0.217 0.274 0.147 0.118 0.098 0.190 0.081 0.198 0.147 0.160 0.127 0.186 0.133 0.095 0.167
PSNR 15.427 8.386 16.189 15.365 22.240 7.930 21.433 25.158 14.965 21.129 24.063 19.481 10.203 25.891 26.146 18.267
FreeSplatter SSIM 0.381 0.566 0.399 0.577 0.822 0.658 0.652 0.851 0.586 0.732 0.739 0.718 0.548 0.776 0.827 0.655
LPIPS 0.434 0.304 0.379 0.238 0.139 0.246 0.233 0.111 0.311 0.228 0.207 0.211 0.322 0.169 0.132 0.244
PSNR 13.207 15.304 15.559 17.032 19.373 13.684 20.432 23.233 17.266 19.399 20.146 21.977 13.602 22.305 22.993 18.368
VGGT SSIM 0.511 0.735 0.619 0.744 0.868 0.827 0.744 0.898 0.754 0.807 0.803 0.836 0.723 0.832 0.860 0.771
LPIPS 0.404 0.204 0.282 0.170 0.130 0.173 0.206 0.100 0.214 0.176 0.194 0.164 0.238 0.146 0.107 0.194
PSNR 24.615 22.975 26.135 26.706 28.720 29.818 27.434 33.895 26.526 25.759 31.803 27.280 28.393 34.096 35.101 28.617
Ours(wo Opt.) SSIM 0.751 0.822 0.763 0.874 0.939 0.955 0.887 0.964 0.879 0.913 0.922 0.917 0.884 0.932 0.951 0.890
LPIPS 0.312 0.153 0.201 0.119 0.082 0.065 0.141 0.060 0.153 0.128 0.143 0.122 0.144 0.091 0.068 0.132
PSNR 25.909 25.376 27.419 26.491 33.835 32.678 31.306 28.887 28.919 27.382 33.239 32.483 30.114 34.646 35.083 30.251
Ours SSIM 0.802 0.860 0.800 0.872 0.988 0.961 0.929 0.945 0.910 0.899 0.925 0.952 0.902 0.937 0.949 0.909
LPIPS 0.231 0.133 0.173 0.104 0.015 0.060 0.100 0.090 0.121 0.149 0.117 0.080 0.124 0.082 0.067 0.110
Table S1: Detailed NVS experiments on DTU datasets.
Scan Office0 Office1 Office2 Room0 Room1 Mean
PSNR 24.503 29.560 18.541 18.452 20.942 22.400
3DGS-GT SSIM 0.848 0.883 0.749 0.786 0.755 0.804
LPIPS 0.417 0.348 0.456 0.428 0.393 0.409
PSNR 24.769 27.655 17.313 18.677 18.372 21.357
2DGS-GT SSIM 0.852 0.877 0.740 0.788 0.742 0.800
LPIPS 0.416 0.353 0.469 0.425 0.417 0.416
PSNR 24.867 28.187 18.165 18.630 20.535 22.077
PGSR-GT SSIM 0.793 0.880 0.701 0.739 0.753 0.773
LPIPS 0.450 0.293 0.494 0.457 0.398 0.418
PSNR 25.736 28.701 17.818 18.287 20.870 22.282
FatesGS-GT SSIM 0.863 0.875 0.747 0.783 0.761 0.806
LPIPS 0.400 0.337 0.461 0.426 0.388 0.402
PSNR 25.428 27.012 17.730 18.772 21.346 22.058
3DGS SSIM 0.861 0.861 0.745 0.785 0.758 0.802
LPIPS 0.411 0.357 0.464 0.434 0.391 0.411
PSNR 23.905 26.530 17.560 18.530 19.844 21.274
2DGS SSIM 0.850 0.871 0.742 0.783 0.749 0.799
LPIPS 0.415 0.354 0.462 0.433 0.398 0.412
PSNR 24.280 25.815 17.366 18.774 20.973 21.441
PGSR SSIM 0.780 0.871 0.691 0.743 0.750 0.767
LPIPS 0.465 0.301 0.509 0.463 0.404 0.428
PSNR 24.723 30.608 17.788 19.013 20.544 22.535
FatesGS SSIM 0.856 0.900 0.744 0.789 0.748 0.808
LPIPS 0.408 0.325 0.461 0.422 0.395 0.402
PSNR 22.460 24.403 12.632 15.656 16.582 18.347
CF-3DGS SSIM 0.778 0.734 0.603 0.726 0.636 0.695
LPIPS 0.336 0.314 0.477 0.427 0.382 0.387
PSNR 20.143 20.134 8.848 12.575 11.074 14.555
FreeSplatter SSIM 0.742 0.635 0.514 0.694 0.556 0.628
LPIPS 0.422 0.471 0.533 0.409 0.454 0.458
PSNR 16.141 20.113 10.272 9.489 9.219 13.047
VGGT SSIM 0.620 0.630 0.494 0.476 0.436 0.531
LPIPS 0.463 0.385 0.538 0.577 0.606 0.514
PSNR 24.230 28.125 19.985 22.502 22.711 23.511
Ours(wo Opt.) SSIM 0.779 0.773 0.720 0.773 0.743 0.758
LPIPS 0.403 0.368 0.419 0.373 0.349 0.382
PSNR 40.059 41.109 24.113 38.218 35.461 35.792
Ours SSIM 0.978 0.968 0.931 0.974 0.976 0.965
LPIPS 0.073 0.137 0.131 0.079 0.057 0.095
Table S2: Detailed NVS experiments on Replica datasets.
Scan 5a48d4 5a618c 5a69c4 5b908d 59338e Mean
PSNR 18.990 16.663 17.389 18.280 17.341 17.733
3DGS-GT SSIM 0.696 0.469 0.471 0.518 0.582 0.547
LPIPS 0.411 0.516 0.516 0.572 0.480 0.499
PSNR 17.751 17.420 16.283 17.722 17.027 17.241
2DGS-GT SSIM 0.687 0.469 0.462 0.520 0.576 0.543
LPIPS 0.418 0.512 0.522 0.570 0.483 0.501
PSNR 18.755 17.003 16.273 19.026 17.098 17.631
PGSR-GT SSIM 0.672 0.461 0.483 0.536 0.567 0.544
LPIPS 0.382 0.529 0.508 0.545 0.486 0.490
PSNR 16.678 16.324 16.423 19.991 13.396 16.563
FatesGS-GT SSIM 0.678 0.472 0.459 0.524 0.542 0.535
LPIPS 0.416 0.511 0.523 0.558 0.521 0.506
PSNR 24.075 16.587 18.005 18.310 16.788 18.753
3DGS SSIM 0.787 0.465 0.476 0.519 0.578 0.565
LPIPS 0.326 0.517 0.510 0.574 0.486 0.483
PSNR 20.773 17.093 16.617 18.195 16.371 17.810
2DGS SSIM 0.722 0.464 0.467 0.520 0.573 0.549
LPIPS 0.386 0.513 0.519 0.569 0.489 0.495
PSNR 22.861 16.725 17.730 19.215 16.742 18.655
PGSR SSIM 0.756 0.461 0.501 0.540 0.558 0.563
LPIPS 0.309 0.528 0.497 0.548 0.492 0.475
PSNR 20.626 14.752 18.949 19.419 17.816 18.312
FatesGS SSIM 0.772 0.418 0.498 0.526 0.612 0.565
LPIPS 0.333 0.537 0.496 0.552 0.466 0.477
PSNR 9.222 14.539 12.652 12.736 15.536 12.937
CF-3DGS SSIM 0.509 0.387 0.398 0.334 0.533 0.432
LPIPS 0.451 0.454 0.465 0.517 0.378 0.453
PSNR 16.111 16.398 16.396 14.015 13.026 15.189
FreeSplatter SSIM 0.629 0.415 0.451 0.493 0.452 0.488
LPIPS 0.390 0.503 0.508 0.521 0.533 0.491
PSNR 11.565 15.799 12.832 14.345 12.493 13.407
VGGT SSIM 0.590 0.543 0.478 0.478 0.546 0.527
LPIPS 0.403 0.347 0.425 0.457 0.425 0.411
PSNR 17.721 21.681 17.158 21.751 19.262 19.515
Ours(wo Opt.) SSIM 0.700 0.670 0.551 0.604 0.644 0.634
LPIPS 0.325 0.268 0.383 0.424 0.362 0.353
PSNR 28.986 32.801 29.437 33.096 29.362 30.737
Ours SSIM 0.977 0.973 0.951 0.964 0.957 0.965
LPIPS 0.038 0.034 0.065 0.055 0.068 0.052
Table S3: Detailed NVS experiments on BlendedMVS datasets.
B.2 Detailed optimization speed experiments
Table 5 presents the average time consumption of optimization-based reconstruction algorithms on Replica datasets, while Tables S4 provide the detailed per-scene evaluation results.
3DGS 2DGS FatesGS PGSR FSFSplatter
Office0 407.72 800.88 638.38 873.45 105.19
Office1 383.28 938.22 975.06 1102.7 103.22
Office2 438.57 825.10 811.89 1037.2 112.44
Room0 434.96 775.38 838.49 853.54 106.88
Room1 482.75 862.31 1315.8 1224.1 109.23
Average 429.45 840.37 915.93 1018.1 107.39
Table S4: Detailed optimization speed experiments on Replica datasets.
B.3 Detailed pose estimation experiments
Table 4 presents the average ATE RMSE error on DTU and Replica datasets, while Tables S5 provide the detailed per-scene evaluation results.
CF-3DGS FreeSplatter Fast3R Dust3r VGGT Ours
Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans
Replica Office0 74.930 38.575 51.045 28.432 4.236 4.417 35.950 3.916 0.794 2.326 0.800 2.319
Office1 43.651 31.719 62.805 42.110 11.592 4.145 60.846 3.468 0.778 3.625 0.740 3.426
Office2 24.612 190.52 44.780 128.35 0.526 4.870 29.353 12.435 0.442 3.156 0.419 2.980
Room0 44.478 173.42 32.990 229.53 0.497 2.343 38.550 13.274 0.889 8.321 0.893 8.308
Room1 38.663 243.30 44.063 109.78 0.306 2.411 34.774 5.936 0.332 1.760 0.318 1.695
DTU scan24 23.123 10.736 4.830 30.796 1.501 3.564 29.952 1.880 0.147 0.383 0.152 0.388
scan37 56.389 95.731 10.628 8.832 5.295 18.280 47.159 15.749 0.461 1.446 0.441 1.458
scan40 50.555 70.607 6.443 15.590 1.289 5.681 59.713 7.021 0.626 1.418 0.628 1.438
scan55 30.339 69.516 4.707 15.411 1.325 11.293 49.758 4.359 0.203 1.209 0.216 1.221
scan63 39.299 79.984 3.520 30.415 4.760 17.137 37.742 17.378 0.530 2.206 0.516 2.203
scan65 11.181 94.896 6.210 39.433 2.124 16.450 25.381 12.917 0.230 1.241 0.276 1.125
scan69 54.851 72.511 2.941 13.699 1.896 6.562 41.689 7.427 0.128 1.853 0.092 1.826
scan83 42.331 86.062 1.877 12.644 6.083 36.981 77.169 26.918 0.325 2.530 0.487 2.181
scan97 29.318 67.483 16.416 30.376 2.506 12.922 56.655 13.913 0.308 1.073 0.313 1.069
scan105 54.325 67.535 3.267 26.114 3.190 26.857 34.217 8.835 0.219 2.054 0.210 2.108
scan106 54.349 72.003 5.200 8.806 0.460 10.196 41.294 4.361 0.359 1.196 0.366 1.215
scan110 40.470 61.468 10.383 13.227 1.909 7.209 42.506 8.192 0.296 0.810 0.301 0.761
scan114 8.837 61.963 5.450 38.759 1.854 10.781 26.354 10.068 0.357 1.123 0.363 1.086
scan118 43.873 81.839 2.640 3.735 2.032 15.600 35.484 10.656 0.081 0.576 0.089 0.502
scan122 34.831 71.341 4.121 23.594 2.499 13.661 39.012 8.106 0.244 0.366 0.257 0.258
Table S5: Detailed pose estimation experiments on Replica and DTU datasets.